# Performance Comparison of Gate-Based and Adiabatic Quantum Computing for Power Flow Analysis

### v1

@@ Abstract
In this paper, we present the first direct comparison between gate-based quantum computing (GQC) and adiabatic quantum computing (AQC) for solving the AC power flow (PF) equations. Building on the Adiabatic Quantum Power Flow (AQPF) algorithm originally designed for annealing platforms, we adapt it to the Quantum Approximate Optimization Algorithm (QAOA). The PF equations are reformulated as a combinatorial optimization problem. Numerical experiments on a 4-bus test system assess solution accuracy and computational time. Results from QAOA are benchmarked against those obtained using D-Wave’s Advantage™ system and Fujitsu’s latest generation Digital Annealer, i.e., Quantum-Inspired Integrated Optimization software (QIIO). The findings provide quantitative insights into the performance trade-offs, scalability, and practical viability of GQC versus AQC paradigms for PF analysis, highlighting the potential of quantum algorithms to address the computational challenges associated with modern electricity networks in the Noisy Intermediate-Scale Quantum (NISQ).
@@
This work is part of the DATALESs project (project number 482.20.602), jointly financed by Netherlands Organization for Scientific Research (NWO) and National Natural Science Foundation of China (NSFC). The work utilized the Dutch national e-infrastructure with the support of the SURF Cooperative (grant number EINF-6569). The authors gratefully acknowledge Heinz Wilkening and the service contract ”Quantum Computing for Load Flow” (contract number 690523) with the European Commission Directorate-General Joint Research Centre (EC DG JRC). Furthermore, the authors acknowledge TNO for the access to TNO’s Quantum Application Lab Facility, with special thanks to Frank Phillipson for his assistance. The authors also thank Fujitsu Technology Solutions for granting access to the QIIO software, with special appreciation to Matthieu Parizy for his help.
I Introduction
@@
Power flow (PF) analysis is a foundational task in electricity networks, used to compute the complex voltages at all buses given specified loads, generation, and network topology. These voltages determine the active/reactive power as well as the current flows on each line, thereby underpinning system operation and planning [1]. In alternating-current (AC) networks, PF analysis is governed by Kirchhoff’s laws and leads to a set of non-linear, non-convex equations. In practice, since the AC PF equations cannot be solved analytically, practitioners rely on iterative numerical methods, e.g., Gauss–Seidel (GS) or Newton–Raphson (NR) to find steady-state voltage solutions [2]. While these classical solvers usually succeed, they can fail in large-scale or ill-conditioned cases. For example, GS depends heavily on the initial guess and often diverges for certain operating conditions, whereas NR may not converge if the Jacobian matrix becomes singular. NR is also computationally expensive and can perform poorly in cases with heavy loading or high renewable penetration [3]. In modern electricity networks with many distributed energy resources, these convergence failures can undermine reliability and lead to erroneous solutions. Thus, there is an urgent need for PF algorithms that are both computationally efficient and numerically robust for modern electricity networks [4, 5].
@@
To address the aforementioned challenge, a fundamentally different approach is to reformulate PF as a combinatorial optimization problem. One can discretize bus complex voltages using spin/binary decision variables, thus transforming the PF equations into an Ising model and/or quadratic unconstrained binary optimization (QUBO) representation (e.g., [6]). This combinatorial reformulation allows for the use of optimization solvers, but it also renders the problem strongly NP-hard [7]. As the electricity network grows in size, the number of spin/binary decision variables increases rapidly, e.g., tens per bus, so exhaustive search becomes intractable. In addition, combinatorial reformulation introduces a higher-order polynomial that must be quadratized, further increasing complexity [8]. Therefore, while combinatorial PF analysis offers a new viewpoint, it poses a large-scale NP-hard optimization problem with rapidly growing dimension.
@@
Quantum computing has recently emerged as a promising paradigm for tackling combinatorial optimization problems, among others. By harnessing quantum superposition and entanglement, quantum algorithms can explore exponentially large solution spaces more efficiently than classical approaches (e.g., [9]). In other words, as the problem size increases, classical solvers tend to plateau, while quantum hardware, if idealized, could maintain an advantage. However, current devices are in practice limited [10]. Contemporary quantum devices are divided into two main paradigms: gate-based (circuit model) quantum computing (GQC) and adiabatic (quantum annealing) quantum computing (AQC). GQC hardware uses sequences of quantum gates on qubits in discrete time steps. They are very flexible in algorithm design and, in theory, can offer speedups for certain problems. For example, the Harrow-Hassidim-Lloyd (HHL) algorithm can solve linear equations in sub-exponential time, given enough qubits. However, GQC algorithms tend to require many qubits and deep circuits even for modest problem sizes, which makes them highly susceptible to decoherence and gate errors on present noisy intermediate-scale quantum (NISQ) hardware. In practice, the limited qubit counts and noise levels mean that most GQC experiments are performed on small test problems or on high-fidelity simulators rather than on real quantum hardware [11].
@@
On the other hand, AQC, exemplified by quantum annealers, evolves a quantum system continuously from an easy-to-prepare ground state to the ground state of a problem Hamiltonian. Adiabatic devices, so-called Ising machines, directly implement energy minimization on spin/binary decision variables. They are generally more noise-tolerant because the computation is analog and not gate-based [12]. For example, D-Wave’s Advantage™ system (QA)111www.dwavequantum.com contains on the order of 5,000 superconducting qubits with sparse connectivity, and Fujitsu’s latest generation Digital Annealer (DA), i.e., Quantum-Inspired Integrated Optimization software (QIIO)222en-portal.research.global.fujitsu.com/kozuchi, can emulate annealing on up to 100,000 fully-connected binary variables at room temperature [8]. These Ising machines specialize in solving NP-hard quadratic optimization problems. Nevertheless, they do not guarantee a perfect solution, but often find high-quality minima of the associated problem Hamiltonian [6].
@@
Both paradigms have been applied to combinatorial optimization problems in the literature (e.g., [13, 14]). A popular GQC approach is the Quantum Approximate Optimization Algorithm (QAOA), a hybrid quantum-classical variational algorithm that encodes the cost of a combinatorial optimization problem into a parameterized quantum circuit (PQC) [15, 16]. In theory, QAOA can yield better approximation ratios than classical heuristics for problems, such as Max-Cut or graph partitioning [17]. Nevertheless, to date, QAOA has only been implemented on small problem sizes, usually via simulation. Though quantum/digital annealers have shown promise on optimization benchmarks. For instance, a recent study comparing QAOA and AQC found that the analog annealer outperformed the GQC protocol on available machines [18].
@@
In our prior study [8], we introduced a combinatorial PF analysis by discretizing the AC PF equations into an Ising model and a QUBO representation and solved it using Ising machines. Numerical experiments on small test systems demonstrated that QA and QIIO can indeed recover accurate PF solutions and handle ill-conditioned cases. However, to date, there has been no GQC implementation or comparative study of the proposed combinatorial PF analysis. In particular, it is unclear how QAOA (on a GQC hardware or simulator) would perform relative to annealing-based hardware for PF analysis. To address this gap, the present study implements the combinatorial PF analysis using both paradigms. We implement QAOA using PennyLane’s lightning.qubit statevector simulator333docs.pennylane.ai/projects/lightning/en/stable/lightning_qubit/device.html for a standard 4-bus test system [19], and run the same problem on two Ising machines, i.e., QA and QIIO. We then compare the three solvers in terms of solution accuracy and computational time. The main contributions of this paper are:
  • providing the first implementation of the combinatorial PF analysis using QAOA, and
  • presenting a comprehensive comparison between GQC and AQC for PF analysis in the NISQ era.
II Combinatorial Power Flow Analysis
@@
Here, $N$ is the number of buses; $P_{i}$ and $Q_{i}$ are respectively the net active and reactive power at bus $i$; $P_{i}^{\mathrm{G}}$ and $Q_{i}^{\mathrm{G}}$ are respectively the generated active and reactive power at bus $i$; $P_{i}^{\mathrm{D}}$ and $Q_{i}^{\mathrm{D}}$ are respectively the consumed active and reactive power at bus $i$. Traditionally, 1 is solved using iterative numerical methods, such as Newton-Raphson (NR), which is susceptible to the aforementioned limitations.
@@
To further prepare 1 for an Ising model, 2 can be re-ordered, as:
@@
Next, the continuous variables $\mu_{i}$ and $\omega_{i}$ from 3 must be discretized. One possible approach is to introduce multiple spin decision variables per bus $i$, each associated with a predefined increment. These spin variables determine whether the increment is added to or subtracted from a given base value for $\mu_{i}$ and $\omega_{i}$. While this method provides a fine-grained search space, it requires a large number of spin decision variables per bus $i$ for both $\mu_{i}$ and $\omega_{i}$, which significantly increases the dimensionality of the combinatorial optimization problem.
@@
To further obtain a combinatorial optimization problem suitable for an Ising model, 1 is recast as the minimization of the sum of squared residuals for all terms:
@@
$\Delta\omega_{i}$ $=\exp\left(\ln(0.05)+\frac{\text{it}\cdot\left(\ln(1\times 10^{-5})-\ln(0.05)\right)}{\text{it}_{\max}}\right).$ (6b)
Here, ‘it’ is the iteration counter and $\text{it}_{\text{max}}$’ is the maximum number of iterations. Since $\mu$ and $\omega$ are expressed in p.u., the maximum and minimum values for $\Delta\mu$ and $\Delta\omega$ are chosen accordingly to remain within numerically stable ranges.
Substituting 4 into 3a yields:
$\begin{split}{}&P_{i}=\sum_{k=1}^{N}\bigg[\mu_{i}^{0}G_{ik}\mu_{k}^{0}+\omega_{i}^{0}G_{ik}\omega_{k}^{0}+\omega_{i}^{0}B_{ik}\mu_{k}^{0}-\mu_{i}^{0}B_{ik}\omega_{k}^{0}\bigg]\\ &\quad+\bigg[\mu_{i}^{0}G_{ik}s_{k}^{\mu}\Delta\mu_{k}+s_{i}^{\mu}\Delta\mu_{i}G_{ik}\mu_{k}^{0}+\omega_{i}^{0}G_{ik}s_{k}^{\omega}\Delta\omega_{k}\\ &\quad+s_{i}^{\omega}\Delta\omega_{i}G_{ik}\omega_{k}^{0}+\omega_{i}^{0}B_{ik}s_{k}^{\mu}\Delta\mu_{k}+s_{i}^{\omega}\Delta\omega_{i}B_{ik}\mu_{k}^{0}\\ &\quad-\mu_{i}^{0}B_{ik}s_{k}^{\omega}\Delta\omega_{k}-s_{i}^{\mu}\Delta\mu_{i}B_{ik}\omega_{k}^{0}\bigg]\\ &\quad+\bigg[s_{i}^{\mu}\Delta\mu_{i}G_{ik}s_{k}^{\mu}\Delta\mu_{k}+s_{i}^{\omega}\Delta\omega_{i}G_{ik}s_{k}^{\omega}\Delta\omega_{k}\\ &\quad+s_{i}^{\omega}\Delta\omega_{i}B_{ik}s_{k}^{\mu}\Delta\mu_{k}-s_{i}^{\mu}\Delta\mu_{i}B_{ik}s_{k}^{\omega}\Delta\omega_{k}\bigg].\end{split}$ (7)
Here, the three bracketed expressions correspond, respectively, to the constant, linear, and quadratic contributions to $P_{i}$.
Similarly, substituting 4 into 3b yields:
$\begin{split}{}&Q_{i}=\sum_{k=1}^{N}\bigg[\omega_{i}^{0}G_{ik}\mu_{k}^{0}-\mu_{i}^{0}G_{ik}\omega_{k}^{0}-\mu_{i}^{0}B_{ik}\mu_{k}^{0}-\omega_{i}^{0}B_{ik}\omega_{k}^{0}\bigg]\\ &\quad+\bigg[\omega_{i}^{0}G_{ik}s_{k}^{\mu}\Delta\mu_{k}+s_{i}^{\omega}\Delta\omega_{i}G_{ik}\mu_{k}^{0}-\mu_{i}^{0}G_{ik}s_{k}^{\omega}\Delta\omega_{k}\\ &\quad-s_{i}^{\mu}\Delta\mu_{i}G_{ik}\omega_{k}^{0}-\mu_{i}^{0}B_{ik}s_{k}^{\mu}\Delta\mu_{k}-s_{i}^{\mu}\Delta\mu_{i}B_{ik}\mu_{k}^{0}\\ &\quad-\omega_{i}^{0}B_{ik}s_{k}^{\omega}\Delta\omega_{k}-s_{i}^{\omega}\Delta\omega_{i}B_{ik}\omega_{k}^{0}\bigg]\\ &\quad+\bigg[s_{i}^{\omega}\Delta\omega_{i}G_{ik}s_{k}^{\mu}\Delta\mu_{k}-s_{i}^{\mu}\Delta\mu_{i}G_{ik}s_{k}^{\omega}\Delta\omega_{k}\\ &\quad-s_{i}^{\mu}\Delta\mu_{i}B_{ik}s_{k}^{\mu}\Delta\mu_{k}-s_{i}^{\omega}\Delta\omega_{i}B_{ik}s_{k}^{\omega}\Delta\omega_{k}\bigg],\end{split}$ (8)
In this study, we develop an Ising model for the proposed combinatorial PF analysis. An Ising model [20] is formulated as the problem of finding the spin decision variable vector $\mathbf{s}\in\{\pm 1\}^{n}$ that minimizes:
@@
Note that 7 and 8 already contain quadratic terms resulting from the interactions between pairs of spin variables, e.g., $s_{i}^{\mu}\Delta\mu_{i}G_{ik}s_{k}^{\mu}\Delta\mu_{k}$ in 7. Therefore, substituting 7 and 8 into 5 yields a fourth-order polynomial in the spin variables. To solve this minimization problem, we use two quantum computing paradigms: gate-based quantum computing (GQC) and adiabatic quantum computing (AQC). For GQC, we employ the Quantum Approximate Optimization Algorithm (QAOA), which utilizes a parameterized quantum circuit (PQC) and variational optimization, thereby allowing direct encoding and optimization of problems with higher-order interactions. For AQC, we use two Ising machines: D-Wave’s Advantage™ system (QA) and Fujitsu’s Quantum-Inspired Integrated Optimization software (QIIO). Since current Ising machines can only handle quadratic interactions, we use the Python package PyQUBO [^4] https://pyqubo.readthedocs.io to reduce higher-order interactions to quadratic ones. QIIO natively supports higher-order reduction, thus allowing for the direct implementation of the fourth-order polynomial.
The iterative scheme used for the combinatorial PF analysis is outlined in Algorithm 1. First, the generation and demand power vectors, $\mathbf{P}^{\mathrm{G}}$, $\mathbf{P}^{\mathrm{D}}$, and $\mathbf{Q}^{\mathrm{D}}$, along with the admittance matrix $\mathbf{Y}$, are initialized according to the given power system data (lines 1-3). The increment vectors, $\Delta\mu$ and $\Delta\omega$, and the initial real and imaginary voltage vectors, $\mathbf{\mu}^{0}$ and $\mathbf{\omega}^{0}$, are then assigned by user-defined values (lines 4-5). Based on these initializations, the corresponding active and reactive power vectors, $\mathbf{P}$ and $\mathbf{Q}$, are computed (line 6), excluding the slack bus entries, $P_{1}$ and $Q_{1}$, to remain consistent with the PF equations. Next, the problem Hamiltonian 5 is evaluated for the initial $\mathbf{\mu}^{0}$ and $\mathbf{\omega}^{0}$ (line 7), providing a first estimate of the solution. A convergence threshold $\epsilon$ is set and the iteration counter ‘it‘ is initialized to zero (line 8). During each iteration, the problem Hamiltonian is minimized with a given solver (line 11). In this study, three different solvers are used, i.e., QAOA, QA, and QIIO. The resulting spin variable vector $\mathbf{s}\in\{\pm 1\}^{2N}$ is used to update the voltage components $\mu_{i}$ and $\omega_{i}$ according to 4 (line 11). If 5 falls below $\epsilon$, the corresponding complex voltages, $\mathbf{\mu}+j\mathbf{\omega}$, are accepted as the PF solution. Otherwise, the base voltage values are reset, $\mathbf{\mu}^{0}:=\mu$ and $\mathbf{\omega}^{0}:=\omega$ (line 14), and the increments are adjusted using 6. The loop continues with the updated Hamiltonian until the algorithm converges to a solution.
@@
6:Compute initial active and reactive power vectors $\mathbf{P}=[P_{2},\dots,P_{N}]$ and $\mathbf{Q}=[Q_{2},\dots,Q_{N}]$ using 3
7:Evaluate 5 with $\mathbf{\mu}^{0}$ and $\mathbf{\omega}^{0}$
@@
10:  Minimize 5 with a given solver
11:  Update voltage vectors $\mathbf{\mu}$ and $\mathbf{\omega}$ using 4
12:  Calculate $\mathbf{P}$ and $\mathbf{Q}$ with updated $\mathbf{\mu}$ and $\mathbf{\omega}$ using 3
13:  Evaluate 5 updated $\mathbf{\mu}$ and $\mathbf{\omega}$
@@
15:  Update increment vectors $\Delta\mu$ and $\Delta\omega$ using 6
@@
III Quantum Approximate Optimization Algorithm

### v2

@@ Abstract
We present the first direct comparison between gate-based quantum computing (GQC) and adiabatic quantum computing (AQC) paradigms for solving the AC power flow (PF) equations. The PF problem is reformulated as a combinatorial optimization problem. For the GQC approach, the Quantum Approximate Optimization Algorithm (QAOA) is employed, while for the AQC approach, the problem is formulated as an Ising model. Numerical experiments on a 4-bus test system evaluate solution accuracy and computational performance. Results obtained using QAOA are benchmarked against those produced by D-Wave’s Advantage™ system and Fujitsu’s latest-generation Digital Annealer, implemented through the Quantum-Inspired Integrated Optimization (QIIO) software. The findings provide quantitative insights into the performance trade-offs, scalability, and practical viability of GQC and AQC paradigms for PF analysis, highlighting the potential of quantum optimization algorithms to address the computational challenges associated with the operation of modern electricity grids in the fault-tolerant era.
keywords:
@@
{highlights}
First direct comparison of gate-based and adiabatic quantum computing for PF analysis.
Formulation of the AC PF problem as a combinatorial optimization problem.
Implementation of QAOA on gate-based simulators and annealers on Ising machines.
Evaluation of solutions and computational performance on a standard test system.
Insights into consistency of quantum approaches with classical PF solvers.
1 Introduction
@@
Power flow (PF) analysis is a foundational task in electricity grids, used to compute the complex voltages at all buses given specified loads, generation, and grid topology. These voltages determine the nodal active and reactive power as well as the current flows on each line, thereby underpinning system operation and planning Hadi Saadat (1999); Xiao and Li (2024). In alternating-current (AC) grids, PF equations are governed by Kirchhoff’s laws and lead to a set of nonlinear, nonconvex equations. In practice, since the AC PF equations cannot be solved analytically, practitioners rely on iterative numerical methods, e.g., Gauss-Seidel (GS) or Newton-Raphson (NR) to find steady-state voltage solutions Arrillaga and Smith (1998). While these classical solvers usually succeed, they can fail in large-scale or ill-conditioned cases. For example, GS depends heavily on the initial guess and often diverges under certain operating conditions, whereas NR may fail to converge if the Jacobian matrix becomes singular. NR is also computationally intensive and can perform poorly under heavy loading or high renewable penetration Mokryani et al. (2016). In modern electricity grids with many distributed energy resources, such convergence failures can undermine reliability and lead to inaccurate solutions. Thus, there is an increasing need for algorithms for the PF problem that are both computationally efficient and numerically robust to account for the challenges of modern electricity grids Tripathy et al. (1982); Tostado-Véliz et al. (2021).
@@
To address the aforementioned challenge, a fundamentally different approach is to reformulate the PF problem as a combinatorial optimization problem by discretizing bus complex voltages using spin/binary decision variables, thus transforming the AC PF equations into an Ising model and/or quadratic unconstrained binary optimization (QUBO) representation, e.g., Kaseb et al. (2025a). This combinatorial reformulation allows for the use of optimization solvers, but it also renders the problem strongly NP-hard Bienstock and Verma (2019). As the electricity grid grows, the number of decision variables increases rapidly, making exhaustive search intractable. In addition, combinatorial reformulation introduces a higher-order polynomial that must be quadratized, further increasing the complexity Kaseb et al. (2024b). Thus, while combinatorial PF reformulation offers a new perspective, it poses a large-scale, NP-hard optimization problem with rapidly growing dimensions.
@@
Quantum computing has recently emerged as a promising paradigm for tackling combinatorial optimization problems, among others, e.g., Ganesan et al. (2021); Ali et al. (2025). By harnessing quantum superposition and entanglement, quantum algorithms can potentially explore exponentially large solution spaces more efficiently than classical approaches, e.g., Au-Yeung et al. (2023). In other words, as the problem size increases, classical solvers tend to plateau, while quantum hardware, if idealized, could maintain an advantage. However, current devices are in practice limited Pareek et al. . Contemporary quantum devices are divided into two main paradigms: gate-based (circuit model) quantum computing (GQC) and adiabatic (quantum annealing) quantum computing (AQC). GQC hardware uses sequences of quantum gates on qubits in discrete time steps. They are very flexible in algorithm design and, in theory, can offer speedups for certain problems. For example, the Harrow-Hassidim-Lloyd (HHL) algorithm can solve linear equations in subexponential time, provided enough qubits are available. However, GQC algorithms tend to require many qubits and deep circuits even for modest problem sizes, which makes them highly susceptible to decoherence and gate errors on present noisy intermediate-scale quantum (NISQ) hardware.
@@
In practice, the limited qubit counts and noise levels mean that most GQC experiments are performed on small test problems or on high-fidelity simulators rather than on real quantum hardware Imre (2014); Kaseb et al. (2024a). On the other hand, AQC, exemplified by quantum annealers, evolves a quantum system continuously from an easy-to-prepare ground state to the ground state of a problem Hamiltonian. Adiabatic devices, so-called Ising machines, directly implement energy minimization on spin/binary decision variables. They are generally more noise-tolerant because the computation is analog and not gate-based Kaseb et al. (2025b). For example, D-Wave’s Advantage™ system (QA)111www.dwavequantum.com contains on the order of 5,000 superconducting qubits with sparse connectivity. Fujitsu’s latest generation Digital Annealer (DA), i.e., Quantum-Inspired Integrated Optimization software (QIIO)222en-portal.research.global.fujitsu.com/kozuchi, can emulate annealing on up to 100,000 fully-connected binary variables at room temperature Kaseb et al. (2024b). These Ising machines specialize in solving NP-hard quadratic optimization problems. Nevertheless, they do not guarantee a perfect solution, but often find high-quality minima of the associated problem Hamiltonian Kaseb et al. (2025a).
@@
Both paradigms have been applied to combinatorial optimization problems in the literature, e.g., Dupont et al. (2023); Armas et al. (2024). A popular GQC approach is the Quantum Approximate Optimization Algorithm (QAOA), a hybrid quantum-classical variational algorithm that encodes the cost of a combinatorial optimization problem into a parameterized quantum circuit (PQC) Pelofske et al. (2023); Basso et al. (2022). In theory, QAOA can yield better approximation ratios than classical heuristics for problems, such as Max-Cut or graph partitioning Jing et al. (2023). Nevertheless, to date, QAOA has only been implemented on small problem sizes, usually via simulation. Though quantum/digital annealers have shown promise on optimization benchmarks. For instance, a recent study comparing QAOA and AQC found that the analog annealer outperformed the GQC protocol on available machines Pelofske et al. (2024). Fig.˜1 illustrates the process of solving combinatorial optimization problems using GQC and AQC devices. The workflow comprises problem encoding, algorithm implementation, and hardware execution. Problem encoding refers to formulating the target optimization problem as an Ising model. Algorithm implementation includes quantum circuit design for GQC, and embedding for AQC. Hardware execution involves applying microwave pulse sequences to GQC hardware and realizing minor embedded via programmable couplers on AQC hardware. The figure illustrates IBM hardware for the former, and QA and DA for the latter.
Figure 1: Workflow for solving combinatorial optimization problems using GQC and AQC devices. The process consists of three stages: problem encoding, algorithm implementation, and hardware execution. In the encoding stage, the target optimization problem is formulated as an Ising model. During algorithm implementation, the problem is mapped to quantum circuits for GQC or embedded onto the hardware graph for AQC. In the execution stage, microwave pulse sequences are applied to gate-based quantum processors, e.g., IBM hardware, while minor-embedded Ising models are realized through programmable couplers on devices such as QA and DA.
@@
In our prior study Kaseb et al. (2024b), we introduced a combinatorial PF analysis by discretizing the AC PF equations into an Ising model and a QUBO representation and solved it using Ising machines. Numerical experiments on small test systems demonstrated that QA and QIIO can indeed recover accurate PF solutions and handle ill-conditioned cases. However, to date, there has been no GQC implementation or comparative study of the proposed combinatorial PF problem. In particular, it is unclear how QAOA (on a GQC hardware or simulator) would perform relative to annealing-based hardware for the PF problem. To address this gap, the present study employs both paradigms to solve AC PF equations. We implement QAOA using PennyLane’s lightning.qubit statevector simulator333docs.pennylane.ai/projects/lightning/en/stable/lightning_qubit/device.html for a standard 4-bus test system Grainger and Stevenson Jr. (1994), and run the same problem on two Ising machines, i.e., QA and QIIO. We then compare the three solvers in terms of solution accuracy and computational time. The main contributions of this paper are: (i) providing the first implementation of the combinatorial PF problem using QAOA, and (ii) presenting a comprehensive comparison between GQC and AQC for the AC PF problem in the NISQ era.
2 Combinatorial Power Flow Analysis
@@
Here, $N$ is the number of buses; $P_{i}$ and $Q_{i}$ are respectively the net active and reactive power at bus $i$; $P_{i}^{\mathrm{G}}$ and $Q_{i}^{\mathrm{G}}$ are respectively the generated active and reactive power at bus $i$; $P_{i}^{\mathrm{D}}$ and $Q_{i}^{\mathrm{D}}$ are respectively the consumed active and reactive power at bus $i$. Traditionally, ˜1 is solved using iterative numerical methods, such as NR, which is susceptible to the aforementioned limitations.
@@
To further prepare ˜1 for an Ising model, ˜2 can be re-ordered, as:
@@
Next, the continuous variables $\mu_{i}$ and $\omega_{i}$ from ˜3 must be discretized. One possible approach is to introduce multiple spin decision variables per bus $i$, each associated with a predefined increment. These spin variables determine whether the increment is added to or subtracted from a given base value for $\mu_{i}$ and $\omega_{i}$. While this method provides a fine-grained search space, it requires a large number of spin decision variables per bus $i$ for both $\mu_{i}$ and $\omega_{i}$, which significantly increases the dimensionality of the combinatorial optimization problem.
@@
To further obtain a combinatorial optimization problem suitable for an Ising model, ˜1 is recast as the minimization of the sum of squared residuals for all terms:
@@
$\Delta\omega_{i}$ $=\exp\left(\ln(0.05)+\frac{\text{it}\cdot\left(\ln(1\times 10^{-5})-\ln(0.05)\right)}{\text{it}_{\max}}\right),$ (6b)
where ‘it’ is the iteration counter, $\text{it}_{\text{max}}$’ is the maximum number of iterations. Since $\mu$ and $\omega$ are expressed in p.u., the maximum and minimum values for $\Delta\mu$ and $\Delta\omega$ are chosen accordingly to remain within numerically stable ranges.
Substituting ˜4 into ˜3a yields:
$\begin{split}{}&P_{i}=\sum_{k=1}^{N}\bigg[\mu_{i}^{0}G_{ik}\mu_{k}^{0}+\omega_{i}^{0}G_{ik}\omega_{k}^{0}+\omega_{i}^{0}B_{ik}\mu_{k}^{0}-\mu_{i}^{0}B_{ik}\omega_{k}^{0}\bigg]\\ &\quad+\bigg[\mu_{i}^{0}G_{ik}s_{k}^{\mu}\Delta\mu_{k}+s_{i}^{\mu}\Delta\mu_{i}G_{ik}\mu_{k}^{0}+\omega_{i}^{0}G_{ik}s_{k}^{\omega}\Delta\omega_{k}\\ &\quad+s_{i}^{\omega}\Delta\omega_{i}G_{ik}\omega_{k}^{0}+\omega_{i}^{0}B_{ik}s_{k}^{\mu}\Delta\mu_{k}+s_{i}^{\omega}\Delta\omega_{i}B_{ik}\mu_{k}^{0}\\ &\quad-\mu_{i}^{0}B_{ik}s_{k}^{\omega}\Delta\omega_{k}-s_{i}^{\mu}\Delta\mu_{i}B_{ik}\omega_{k}^{0}\bigg]\\ &\quad+\bigg[s_{i}^{\mu}\Delta\mu_{i}G_{ik}s_{k}^{\mu}\Delta\mu_{k}+s_{i}^{\omega}\Delta\omega_{i}G_{ik}s_{k}^{\omega}\Delta\omega_{k}\\ &\quad+s_{i}^{\omega}\Delta\omega_{i}B_{ik}s_{k}^{\mu}\Delta\mu_{k}-s_{i}^{\mu}\Delta\mu_{i}B_{ik}s_{k}^{\omega}\Delta\omega_{k}\bigg],\end{split}$ (7)
where the three bracketed expressions correspond, respectively, to constant, linear, and quadratic contributions to $P_{i}$.
Similarly, substituting ˜4 into ˜3b yields:
$\begin{split}{}&Q_{i}=\sum_{k=1}^{N}\bigg[\omega_{i}^{0}G_{ik}\mu_{k}^{0}-\mu_{i}^{0}G_{ik}\omega_{k}^{0}-\mu_{i}^{0}B_{ik}\mu_{k}^{0}-\omega_{i}^{0}B_{ik}\omega_{k}^{0}\bigg]\\ &\quad+\bigg[\omega_{i}^{0}G_{ik}s_{k}^{\mu}\Delta\mu_{k}+s_{i}^{\omega}\Delta\omega_{i}G_{ik}\mu_{k}^{0}-\mu_{i}^{0}G_{ik}s_{k}^{\omega}\Delta\omega_{k}\\ &\quad-s_{i}^{\mu}\Delta\mu_{i}G_{ik}\omega_{k}^{0}-\mu_{i}^{0}B_{ik}s_{k}^{\mu}\Delta\mu_{k}-s_{i}^{\mu}\Delta\mu_{i}B_{ik}\mu_{k}^{0}\\ &\quad-\omega_{i}^{0}B_{ik}s_{k}^{\omega}\Delta\omega_{k}-s_{i}^{\omega}\Delta\omega_{i}B_{ik}\omega_{k}^{0}\bigg]\\ &\quad+\bigg[s_{i}^{\omega}\Delta\omega_{i}G_{ik}s_{k}^{\mu}\Delta\mu_{k}-s_{i}^{\mu}\Delta\mu_{i}G_{ik}s_{k}^{\omega}\Delta\omega_{k}\\ &\quad-s_{i}^{\mu}\Delta\mu_{i}B_{ik}s_{k}^{\mu}\Delta\mu_{k}-s_{i}^{\omega}\Delta\omega_{i}B_{ik}s_{k}^{\omega}\Delta\omega_{k}\bigg].\end{split}$ (8)
In this study, we develop an Ising model for the proposed combinatorial PF analysis. An Ising model Goto et al. (2019) is formulated as the problem of finding the spin decision variable vector $\mathbf{s}\in\{\pm 1\}^{n}$ that minimizes:
@@
Note that ˜7 and 8 already contain quadratic terms resulting from the interactions between pairs of spin variables, e.g., $s_{i}^{\mu}\Delta\mu_{i}G_{ik}s_{k}^{\mu}\Delta\mu_{k}$ in ˜7. Therefore, substituting ˜7 and 8 into ˜5 yields a fourth-order polynomial in the spin variables. To solve this minimization problem, we use two quantum computing paradigms: GQC and AQC. For GQC, we employ QAOA, which utilizes a PQC and variational optimization, thereby allowing direct encoding and optimization of problems with higher-order interactions. For AQC, we use two Ising machines: QA and OIIO. Since current Ising machines can only handle quadratic interactions, we use the Python package PyQUBO [^4] https://pyqubo.readthedocs.io to reduce higher-order interactions to quadratic ones. QIIO natively supports higher-order reduction, enabling the direct implementation of the fourth-order polynomial.
The iterative scheme used for the combinatorial PF analysis is outlined in Algorithm˜1. First, the generation and demand power vectors, $\mathbf{P}^{\mathrm{G}}$, $\mathbf{P}^{\mathrm{D}}$, and $\mathbf{Q}^{\mathrm{D}}$, along with the admittance matrix $\mathbf{Y}$, are initialized according to the given power system data (lines 1-3). The increment vectors, $\Delta\mu$ and $\Delta\omega$, and the initial real and imaginary voltage vectors, $\mathbf{\mu}^{0}$ and $\mathbf{\omega}^{0}$, are then assigned by user-defined values (lines 4-5). Based on these initializations, the corresponding active and reactive power vectors, $\mathbf{P}$ and $\mathbf{Q}$, are computed (line 6), excluding the slack bus entries, $P_{1}$ and $Q_{1}$, to remain consistent with the PF equations. Next, the problem Hamiltonian ˜5 is evaluated for the initial $\mathbf{\mu}^{0}$ and $\mathbf{\omega}^{0}$ (line 7), providing a first estimate of the solution. A convergence threshold $\epsilon$ is set and the iteration counter ‘it‘ is initialized to zero (line 8). During each iteration, the problem Hamiltonian is minimized with a given solver (line 11). In this study, three solvers are used: QAOA, QA, and QIIO. The resulting spin variable vector $\mathbf{s}\in\{\pm 1\}^{2N}$ is used to update the voltage components $\mu_{i}$ and $\omega_{i}$ according to ˜4 (line 11). If ˜5 falls below $\epsilon$, the corresponding complex voltages, $\mathbf{\mu}+j\mathbf{\omega}$, are accepted as the PF solution. Otherwise, the base voltage values are reset, $\mathbf{\mu}^{0}:=\mu$ and $\mathbf{\omega}^{0}:=\omega$ (line 14), and the increments are adjusted using ˜6. The loop continues with the updated Hamiltonian until the algorithm converges to a solution.
@@
6:Compute initial active and reactive power vectors $\mathbf{P}=[P_{2},\dots,P_{N}]$ and $\mathbf{Q}=[Q_{2},\dots,Q_{N}]$ using ˜3
7:Evaluate ˜5 with $\mathbf{\mu}^{0}$ and $\mathbf{\omega}^{0}$
@@
10:  Minimize ˜5 with a given solver
11:  Update voltage vectors $\mathbf{\mu}$ and $\mathbf{\omega}$ using ˜4
12:  Calculate $\mathbf{P}$ and $\mathbf{Q}$ with updated $\mathbf{\mu}$ and $\mathbf{\omega}$ using ˜3
13:  Evaluate ˜5 updated $\mathbf{\mu}$ and $\mathbf{\omega}$
@@
15:  Update increment vectors $\Delta\mu$ and $\Delta\omega$ using ˜6
@@
3 Quantum Approximate Optimization Algorithm